OpenStudy (anonymous):

Find the prime factorization of 576. Please show steps

OpenStudy (zarkon):

576=192*3=64*3*3=8*8*3*3=2*2*2*2*2*2*3*3

OpenStudy (zarkon):

\[2^6\cdot3^2\]

OpenStudy (anonymous):

one thing i dont get is why you dont divide by 2 first. It says to divide by 2,3 and 5

OpenStudy (anonymous):

You can divide by whatever numbers you like as long as you use integers. I believe Zarkon chose 3 b/c multiplication properties show that you can. 2 would have worked just as well tho. Here are some multiplication properties that you may find helpful: They are in this format: "A number is a factor of ____ if______" 2 if the last digit is divisible by 2 3 if all the digits added together form a number that is divisible by 3 4 if the last 2 digits form a number that is divisible by 4 5 if the last digit is a 5 or a 0 6 if the number satisfies the rules for both 2 AND 3 9 if all the digits added up form a number that is divisible by 9 10 if the last digit is a 0